If(-7,-3) is the midpoint of segment QP and the coordinates of Q are (6,-3) find the coordinates of P.
Answer by checkley77(12844)

The best method for solving these type of graphing problems is to graph the points.
Thus we have an end point (6,-3) the mid point (-7,-3).
This is a verticle line through y=-3.
The distance from [Q] to the mid point = (6+7)=13 units.
Therefore the missing half will also be 13 units long or -7-13=-20.
So the other end point [P]is(-20,-3).
Even if you don't use the graphing technique to solve the problem then at least use it to do the proof.